A person has weight of 92 kg and he want to fly in the sky with the help of ballon. Each ballon contain 50 mole of hydrogen gas and 0.5 atm pressure and 27' C temperature . If the density of air at given condition is 1.25 gm/l. Then calculate the number of ballon used to fly in the sky.

Grade:11

1 Answers

Vikas TU
7 years ago
From PV = nRT Volume of each balloon => nRT/P = > 50 * 0.0821 * 300/0.5 = > 2,463 litre.
mass of balloon = density of air * volume = > 1.25 * 2463 = > 3078.75 g or 3.078 Kg.
No. of balloons = 92/3.078 = > 29.8 = > 30 balloons.